# Deepfake Generation and Detection: A Benchmark and Survey

### v3

@@ 3.2.3 Frequency Domain
Frequency domain-based forgery detection methods transform image time-domain information into the frequency domain. Works [36, 320, 37, 305, 307] utilize statistical measures of periodic features, frequency components, and frequency characteristic distributions, either globally or in local regions, as evaluation metrics for forgery detection. Specifically, HFI-Net [305] consists of a dual-branch network and four Global-Local Interaction (GLI) modules. It effectively explores multi-level frequency artifacts, obtaining frequency-related forgery clues for face detection. Tan et al[307] introduce a novel frequency-aware approach called FreqNet, which focuses on the high-frequency information of images and combines it with a frequency-domain learning module to learn source-independent features. Furthermore, some approaches combine spatial, temporal, and frequency domains for joint consideration [321, 306], and Guo et al[306] design a spatial-frequency interaction convolution to construct a novel backbone network for Deepfake detection.
@@ 3.2.4 Data Driven
Data-driven forgery detection focuses on learning specific patterns and features from extensive image or video datasets to distinguish between genuine and potentially manipulated images. Some methods [322, 323] believe that images generated by specific models possess unique model fingerprints. Based on this belief, forgery detection can be achieved by focusing on the model’s training. In addition, FakeSpotter [324] introduces the Neuron Coverage Criterion to capture layer-wise neuron activation behavior. It monitors the neural behavior of a deep face recognition system through a binary classifier to detect fake faces. There are also methods [309, 109] that attempt to classify the sources of different components in an image. For instance, Huang et al[109] think that the difference between explicit and implicit identity helps detect face swapping. There are numerous data-driven methods [308, 310, 311, 325], and it is not feasible to discuss each one in detail here.
@@ 3.3.1 Face Super-resolution
$\bullet$ Convolutional Neural Networks. Early works [123, 326, 327] on facial super-resolution based on CNNs aims to leverage the powerful representational capabilities of CNNs to learn the mapping relationship between low-resolution and high-resolution images from training samples. Depending on whether they focus on local details of the image, they can be divided into global methods [123, 122], local methods [326], and mixed methods [327].
$\bullet$ Generative Adversarial Network. GAN aims to achieve the optimal output result through an adversarial process between the generator and the discriminator. This type of method [328, 329] currently dominates the field for flexible and efficient architecture.
@@ 3.3.2 Portrait Style Transfer
$\bullet$ Generative Adversarial Network. The most mature style transfer algorithm is the GAN-based approach [156, 157, 158]. However, due to the relatively poor stability of GANs, it is common for the generated images to contain artifacts and unreasonable components. 3DAvatarGAN [156] bridges the pre-trained 3D-GAN in the source domain with the 2D-GAN trained on an artistic dataset to achieve cross-domain generation. Scenimefy [158] utilizes semantic constraints provided by text models like CLIP to guide StyleGAN generation and applies patch-based contrastive style loss to enhance stylization and fine details further.
$\bullet$ Diffusion-based methods [159, 160, 330] represent the generative process of cross-domain image transfer using diffusion processes. DiffusionGAN3D [330] combines 3DGAN [271] with a diffusion model from text to graphics, introducing relative distance loss and learnable tri-planes for specific scenarios to further enhance cross-domain transformation accuracy.
@@ 3.3.3 Body Animation
$\bullet$ Generative Adversarial Network. GAN-based approaches [147, 331, 146] aim to train a model to generate images whose conditional distribution resembles the target domain, thus transferring information from reference images to target images. CASD [146] is based on a style distribution module using a cross-attention mechanism, facilitating pose transfer between source semantic styles and target poses. VGFlow [331] introduces a visibility-guided flow module to preserve texture and perform style manipulation concurrently. However, existing methods still rely considerably on training samples, and exhibit decreased performance when dealing with actions in rare poses.
$\bullet$ Diffusion-based. The task of body animation using diffusion models aims to utilize diffusion processes to generate the propagation and interaction of movements between body parts based on a reference source. This approach [148, 49, 332] represents a current hot topic in research and implementation. LEO [148] focuses on the spatiotemporal continuity between generated actions, employing the Latent Motion Diffusion Model to represent motion as a series of flow graphs during the generation process. Animate Anyone [332] harnesses the powerful generation capabilities of stable diffusion models combined with attention mechanisms to generate high-quality character animation video.
@@ 3.3.4 Makeup Transfer
$\bullet$ Graphics-based Approaches. Before the introduction of neural networks, traditional computer graphics methods [333, 334] use image gradient editing and physics-based operations to understand the semantics of makeup. By decomposing the input image into multiple layers, each representing different facial information, traditional methods would distort the reference facial image onto the non-makeup target using facial landmarks for each layer. However, due to the limitations of manually designed operators, the output images from traditional methods often appear unnatural, with noticeable artifacts. Additionally, there is a tendency for background information to be modified to some extent.
$\bullet$ Generative Adversarial Network. Early deep learning-based methods [164] aim at fully automatic makeup transfer. However, these methods [163, 335] exhibit poor performance when faced with significant differences in pose and expression between the source and target faces and are unable to handle extreme makeup scenarios well. Some methods [161, 165, 166] proposes their solutions, PSGAN++ [165] comprises the Makeup Distillation Network, Attentive Makeup Morphing module, Style Transfer Network, and Identity Extraction Network, further enhancing the ability of PSGAN [161] to perform targeted makeup transfer with detail preservation. ELeGANt [167], CUMTGAN [170], and HT-ASE [162] explore the preservation of detailed information. ELeGANt [167] encodes facial attributes into pyramid feature maps to retain high-frequency information. Matching facial semantic information, which involves rendering makeup styles onto semantically corresponding positions of a target image, is often overlooked. SSAT [336] introduces the SSCFT module and weakly supervised semantic loss for accurate semantic correspondence. SSAT++ [169] further improves color fidelity matching, but both models are complex with high training complexity. In addition, BeautyREC [168], based on the Transformer with long-range visual dependencies, achieves efficient global makeup transfer with significantly reduced overall model parameters compared to previous works.
@@ 4 Benchmark Results
We evaluate the performance of representative methods for each reviewed field on the most widely used datasets with the data sourced from the original respective papers. Considering the differences in training datasets, testing datasets, and metrics used by different approaches, we strive to compare them fairly in each table to the greatest extent.
@@ 4.1 Main Results on Deepfake Generation
$\bullet$ Results on Face Swapping. Tab. VII displays the performance evaluation results of some representative models on the Face Swapping task using the FF++ [78] dataset. WSC-Swap[204] captures facial region external attribute information and internal identity information through two independent encoders, which achieves good identity preservation and facial pose retention. However, it exhibits sub-optimal performance in facial expression error metrics.
$\bullet$ Results on Face Reenactment. Tab. VIII and Tab. IX show the performance evaluation results on the VoxCeleb [71] dataset for self-reenactment and cross-subject reenactment, respectively. AVFR-GAN [337] achieves better performance by using the multimodal modeling. Tab. X presents quality assessment results on the VoxCeleb2 dataset.
TABLE VII: Results of representative face swapping methods on FF++. Notations: ➊ CelebA-HQ, ➋ FFHQ, ➌ VGGFace, ➍ VGGFace2, ➎ VoxCeleb2.
@@
FaceShifter [182] ➊➋➌ 97.38 2.06 2.96 -
SimSwap [184] 92.83 - 1.53 -
FaceInpainter [186] ➊➋➌ 97.63 - 2.21 -
HifiFace [203] 98.48 - 2.63 -
RAFSwap [11] 96.70 2.92 2.53 -
Xu et al[121] 90.05 2.79 2.46 -
DiffSwap [192] 98.54 5.35 2.45 2.16
FlowFace [189] ➊➋➍ 99.26 - 2.66 -
FlowFace++ [205] ➊➋➍ 99.51 - 2.20 -
StyleIPSB [202] 95.05 2.23 3.58 -
StyleSwap [209] ➌➎ 97.05 5.28 1.56 2.72
WSC-Swap [204] ➊➋➌ 99.88 5.01 1.51 -
TABLE VIII: Results of representative face reenactment methods on VoxCeleb for the self-reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 [226], ➎ MPIIGaze [227], ➏ TalkingHead-1KH. In addition, we use gray to represent data that is partially uncertain.
@@
HyperReenact [16] 0.710 - 0.230 27.10 -
DG [15] 0.831 - - 22.10 0.761
AVFR-GAN [337] - 32.20 - 08.48 0.824
Free-HeadGAN [225] ➊➌➍➎ 0.810 22.16 0.100 35.40 -
HiDe-NeRF [230] ➊➋➐ 0.931 21.90 0.084 - 0.862
TABLE IX: Results of representative face reenactment methods on VoxCeleb for the cross-identity reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 [226], ➎ MPIIGaze [227], ➏ TalkingHead-1KH.
@@
HyperReenact [16] 0.680 - - - -
AVFR-GAN [337] - - - 09.05 -
Free-HeadGAN [225] ➊➌➍➎ 0.789 - - 53.90 13.1
HiDe-NeRF [230] ➊➋➏ 0.786 0.012 0.971 57.00 -
TABLE X: Results of representative face reenactment methods on VoxCeleb2 for quality assessment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ LRW, ➍ CelebV-HQ, ➎ TalkingHead-1KH.
@@
PC-AVS [244] ➋➌ - - 6.880 - 0.886
GC-AVT [245] - - 2.757 - 0.739
Wang et al[97] - 28.92 1.830 - 0.830
PECHead [219] ➋➍➎ 1.590 - 23.05 -
DG [15] 0.721 - - 51.79 0.540
HiDe-NeRF [230] ➊➎ 0.787 - - 61.00 -
TABLE XI: Evaluation results of the models involved in talking face generation on the MEAD dataset. Notations: ➊ MEAD, ➋ LRW, ➌ VoxCeleb2, ➍ HDTF.
@@
Xu et al[247] 0.83 2.36 - 3.500 15.91 30.09/0.850
EMMN [264] ➊➋ - - 2.780/2.870 3.570 - 29.38/0.660
AMIGO [265] ➊➌ - 2.44 2.140/2.440 - 19.59 30.29/0.820
SLIGO [263] 0.88 1.83 - 3.690 - 00000-/0.790
Gan et al[240] ➊➌ - - 2.250/2.470 - 19.69 21.75/0.680
DreamTalk [18] ➊➌➍ - - 2.910/1.930 3.780 - 00000-/0.860
SPACE [266] ➊➌ - - - 3.610 11.68 0-
TalkCLIP [267] - - 3.601/2.415 3.773 - 00000-/0.829
TABLE XII: FID evaluation of different methods. Notations: ➊ FFHQ, ➋ CelebA-HQ, ➌ MM CelebA-HQ, ➍ CelebA, ➎ CelebAText-HQ, ➏ Self-build.
@@
FENeRF [279] GANs+NeRF 12.10
FENeRF [279] GANs+NeRF 28.20
AnyFace [20] GANs 56.75
AnyFace [20] GANs 50.56
TextFace [292] GANs 22.81
TG-3DFace [283] GANs 52.21
TG-3DFace [283] GANs 39.02
HifaFace [277] GANs ➊➋ ➊➋ 04.04
GuidedStyle [291] GANs 41.79
$\bullet$ Results on Talking Face Generation. Tab. XI displays the performance results of various talking face generation approaches on the MEAD [75] dataset since 2023. AMIGO [265] achieves promising results among all methods.
$\bullet$ Results on Facial Attribute Editing. Tab.XII evaluates the quality level of generated images using FID, and Tab.XIII assesses facial reconstruction capabilities using PSNR, LPIPS, and SSIM. Due to different training and testing datasets, quantitative fair comparisons are not possible that just serve as performance demonstrations.

### v4

@@ 3.2.3 Frequency Domain
Frequency domain-based forgery detection methods transform image time-domain information into the frequency domain. Works qian2020thinking ; frank2020leveraging ; li2021frequency ; miao2022hierarchical ; tan2024frequency utilize statistical measures of periodic features, frequency components, and frequency characteristic distributions, either globally or in local regions, as evaluation metrics for forgery detection. Specifically, F3-Net qian2020thinking proposes a dual-branch framework. One frequency-aware branch utilizes Frequency-aware Image Decomposition (FAD) to learn subtle forgery patterns in suspicious images. In contrast, the other branch aims to extract high-level semantics from Local Frequency Statistics (LFS) to describe the frequency-aware statistical differences between real and forged faces. HFI-Net miao2022hierarchical consists of a dual-branch network and four Global-Local Interaction (GLI) modules. It effectively explores multi-level frequency artifacts, obtaining frequency-related forgery clues for face detection. Tan et altan2024frequency introduce a novel frequency-aware approach called FreqNet, which focuses on the high-frequency information of images and combines it with a frequency-domain learning module to learn source-independent features. Furthermore, some approaches combine spatial, temporal, and frequency domains for joint consideration masi2020two ; guo2023constructing , and Guo et alguo2023constructing design a spatial-frequency interaction convolution to construct a novel backbone network for Deepfake detection.
@@ 3.2.4 Data Driven
Data-driven forgery detection focuses on learning specific patterns and features from extensive image or video datasets to distinguish between genuine and potentially manipulated images. Some methods hsu2018learning ; yu2019attributing believe that images generated by specific models possess unique model fingerprints. Based on this belief, forgery detection can be achieved by focusing on the model’s training. In addition, FakeSpotter wang2021fakespotter introduces the Neuron Coverage Criterion to capture layer-wise neuron activation behavior. It monitors the neural behavior of a deep face recognition system through a binary classifier to detect fake faces. There are also methods zhao2021learning ; huang2023implicit that attempt to classify the sources of different components in an image. For instance, Huang et alhuang2023implicit think that the difference between explicit and implicit identity helps detect face swapping. There are numerous data-driven methods dang2020detection ; hu2022finfer ; guo2023hierarchical ; guo2023controllable , and it is not feasible to discuss each one in detail here.
@@ 3.3.1 Face Super-resolution
$\bullet$ Convolutional Neural Networks. Early works chen2019sequential ; hu2019face ; lu2020global on facial super-resolution based on CNNs aims to leverage the powerful representational capabilities of CNNs to learn the mapping relationship between low-resolution and high-resolution images from training samples. Depending on whether they focus on local details of the image, they can be divided into global methods chen2019sequential ; huang2017wavelet , local methods hu2019face , and mixed methods lu2020global .
$\bullet$ Generative Adversarial Network. GAN aims to achieve the optimal output result through an adversarial process between the generator and the discriminator. This type of method zhang2019ranksrgan ; pan2024lpsrgan currently dominates the field for flexible and efficient architecture.
@@ 3.3.2 Portrait Style Transfer
$\bullet$ Generative Adversarial Network. The most mature style transfer algorithm is the GAN-based approach abdal20233davatargan ; xu2022your3demoji ; jiang2023scenimefy . However, due to the relatively poor stability of GANs, it is common for the generated images to contain artifacts and unreasonable components. 3DAvatarGAN abdal20233davatargan bridges the pre-trained 3D-GAN in the source domain with the 2D-GAN trained on an artistic dataset to achieve cross-domain generation. Scenimefy jiang2023scenimefy utilizes semantic constraints provided by text models like CLIP to guide StyleGAN generation and applies patch-based contrastive style loss to enhance stylization and fine details further.
$\bullet$ Diffusion-based methods liu2023portrait ; hur2024expanding ; lei2023diffusiongan3d represent the generative process of cross-domain image transfer using diffusion processes. DiffusionGAN3D lei2023diffusiongan3d combines 3DGAN chan2022efficient with a diffusion model from text to graphics, introducing relative distance loss and learnable tri-planes for specific scenarios to further enhance cross-domain transformation accuracy.
@@ 3.3.3 Body Animation
$\bullet$ Generative Adversarial Network. GAN-based approaches zhao2022thin ; jain2023vgflow ; zhou2022cross aim to train a model to generate images whose conditional distribution resembles the target domain, thus transferring information from reference images to target images. CASD zhou2022cross is based on a style distribution module using a cross-attention mechanism, facilitating pose transfer between source semantic styles and target poses. VGFlow jain2023vgflow introduces a visibility-guided flow module to preserve texture and perform style manipulation concurrently. However, existing methods still rely considerably on training samples, and exhibit decreased performance when dealing with actions in rare poses.
$\bullet$ Diffusion-based. The task of body animation using diffusion models aims to utilize diffusion processes to generate the propagation and interaction of movements between body parts based on a reference source. This approach wang2023leo ; ma2023follow ; hu2023animate represents a current hot topic in research and implementation. LEO wang2023leo focuses on the spatiotemporal continuity between generated actions, employing the Latent Motion Diffusion Model to represent motion as a series of flow graphs during the generation process. Animate Anyone hu2023animate harnesses the powerful generation capabilities of stable diffusion models combined with attention mechanisms to generate high-quality character animation video.
@@ 3.3.4 Makeup Transfer
$\bullet$ Graphics-based Approaches. Before the introduction of neural networks, traditional computer graphics methods guo2009digital ; li2015simulating use image gradient editing and physics-based operations to understand the semantics of makeup. By decomposing the input image into multiple layers, each representing different facial information, traditional methods would distort the reference facial image onto the non-makeup target using facial landmarks for each layer. However, due to the limitations of manually designed operators, the output images from traditional methods often appear unnatural, with noticeable artifacts. Additionally, there is a tendency for background information to be modified to some extent.
$\bullet$ Generative Adversarial Network. Early deep learning-based methods gu2019ladn aim at fully automatic makeup transfer. However, these methods chang2018pairedcyclegan ; chen2019beautyglow exhibit poor performance when faced with significant differences in pose and expression between the source and target faces and are unable to handle extreme makeup scenarios well. Some methods jiang2020psgan ; liu2021psgan++ ; zhong2023sara proposes their solutions, PSGAN++ liu2021psgan++ comprises the Makeup Distillation Network, Attentive Makeup Morphing module, Style Transfer Network, and Identity Extraction Network, further enhancing the ability of PSGAN jiang2020psgan to perform targeted makeup transfer with detail preservation. ELeGANt yang2022elegant , CUMTGAN hao2022cumtgan , and HT-ASE li2023hybrid explore the preservation of detailed information. ELeGANt yang2022elegant encodes facial attributes into pyramid feature maps to retain high-frequency information. Matching facial semantic information, which involves rendering makeup styles onto semantically corresponding positions of a target image, is often overlooked. SSAT sun2022ssat introduces the SSCFT module and weakly supervised semantic loss for accurate semantic correspondence. SSAT++ sun2023ssat further improves color fidelity matching, but both models are complex with high training complexity. In addition, BeautyREC yan2023beautyrec , based on the Transformer with long-range visual dependencies, achieves efficient global makeup transfer with significantly reduced overall model parameters compared to previous works.
@@ 4 Benchmark Results
We firstly introduce the evaluation metrics commonly used for each deepfake tasks, and then we evaluate the performance of representative methods for each reviewed field on the most widely used datasets with the data sourced from the original respective papers. Considering the differences in training datasets, testing datasets, and metrics used by different approaches, we strive to compare them fairly in each table to the greatest extent.
4.1 Metrics
$\bullet$ Face Swapping. The most commonly used objective evaluation metrics for face swapping include ID Ret, Expression Error, Pose Error, and FID. ID Ret is calculated by a pre-trained face recognition model wang2018cosface , measuring the Euclidean distance between the generated face and the source face. A higher ID Ret indicates better preservation of identity information. Expression and pose errors quantify the differences in expression and pose between the generated face and the source face. These metrics are evaluated using a pose estimator ruiz2018fine and a 3D facial model deng2019accurate , extracting expression and pose vectors for the generated and source faces. Lower values for expression error and pose error indicate higher facial expression and pose similarity between the swapped face and the source face. FID  heusel2017gans is used to assess image quality, with lower FID values indicating that the generated images closely resemble authentic facial images in appearance.
$\bullet$ Face Reenactment. Face reenactment commonly uses consistent evaluation metrics, including CSIM, SSIM wang2004image , PSNR, LPIPS zhang2018unreasonable , LMD chen2018lip , and FID. CSIM describes the cosine similarity between the generated and source faces, calculated by ArcFace deng2019arcface , with higher values indicating better performance. SSIM, PSNR, LPIPS, and FID are used to measure the quality of synthesized images. SSIM measures the structural similarity between two images, with higher values indicating a closer resemblance to natural images. PSNR quantifies the ratio between a signal’s maximum possible power and noise’s power, indicating higher quality for higher values. LPIPS assesses reconstruction fidelity using a pre-trained AlexNet krizhevsky2012imagenet to extract feature maps for similarity score computation. As mentioned earlier, FID is used to evaluate image quality. LMD assesses the accuracy of lip shape in generated images or videos, with lower values indicating better model performance.
$\bullet$ Talking Face Generation. Expanding upon face reenactment metrics, talking face generation incorporates additional metrics, including M/F-LMD, Sync, LSE-C, and LSE-D. The LSE-C and LSE-D are usually used to measure lip synchronization effectiveness prajwal2020lip . The landmark distances on the mouth (M-LMD) chen2019hierarchical and the confidence score of SyncNet (Sync) measure synchronization between the generated lip motion and the input audio. F-LMD computes the difference in the average distance of all landmarks between predictions and ground truth (GT) as a measure to assess the generated expression. In addition, there are some meaningful metrics, such as LSE-C and LSE-D for measuring lip synchronization effectiveness prajwal2020lip , AVD li2023one for evaluating identity preservation performance, AUCON ha2020marionette for assessing facial pose and expression jointly, and AGD doukas2023free for evaluating eye gaze changes. These newly proposed evaluation metrics enrich the performance assessment system by targeting various aspects of the model’s performance.
$\bullet$ Facial Attribute Editing. The standard evaluation metrics used in face attribute manipulation are FID, LPIPS zhang2018unreasonable , KID binkowski2018demystifying , PSNR and SSIM. KID is one of the image quality assessment metrics commonly used in face editing work and other generative modeling tasks to quantify the difference in distribution between the generated image and the actual image, with lower KID values indicating better model performance. Some text-guide work will also use the CLIP Score to measure the consistency between the output image and the text, calculated as the cosine similarity between the normalized image and the text embedding. Higher values of CLIP Score indicate better consistency of the generated image with the corresponding text sentence.
@@ 4.2 Main Results on Deepfake Generation
$\bullet$ Results on Face Swapping. Table 7 displays the performance evaluation results of some representative models on the Face Swapping task using the FF++ rossler2019faceforensics++ dataset. WSC-Swapren2023reinforced captures facial region external attribute information and internal identity information through two independent encoders, which achieves good identity preservation and facial pose retention. However, it exhibits sub-optimal performance in facial expression error metrics.
$\bullet$ Results on Face Reenactment. Table 8 and Table 9 show the performance evaluation results on the VoxCeleb nagrani2017voxceleb dataset for self-reenactment and cross-subject reenactment, respectively. AVFR-GAN agarwal2023audio achieves better performance by using the multimodal modeling. 10 presents quality assessment results on the VoxCeleb2 dataset. HiDe-NeRF li2023one represents 3D scenes using canonical appearance fields and implicit deformation fields. It achieves accurate facial attribute modeling by explicitly decoupling facial pose and expression attributes using the deformation module.
Table 7: Results of representative face swapping methods on FF++. Notations: ➊ CelebA-HQ, ➋ FFHQ, ➌ VGGFace, ➍ VGGFace2, ➎ VoxCeleb2.
@@
FaceShifter li2019faceshifter ➊➋➌ 97.38 2.06 2.96 -
SimSwap chen2020simswap 92.83 - 1.53 -
FaceInpainter li2021faceinpainter ➊➋➌ 97.63 - 2.21 -
HifiFace wang2021hififace 98.48 - 2.63 -
RAFSwap xu2022region 96.70 2.92 2.53 -
Xu et alxu2022high 90.05 2.79 2.46 -
DiffSwap zhao2023diffswap 98.54 5.35 2.45 2.16
FlowFace zeng2023flowface ➊➋➍ 99.26 - 2.66 -
FlowFace++ zhang2023flowface++ ➊➋➍ 99.51 - 2.20 -
StyleIPSB jiang2023styleipsb 95.05 2.23 3.58 -
StyleSwap xu2022styleswap ➌➎ 97.05 5.28 1.56 2.72
WSC-Swap ren2023reinforced ➊➋➌ 99.88 5.01 1.51 -
Table 8: Results of representative face reenactment methods on VoxCeleb for the self-reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 kellnhofer2019gaze360 , ➎ MPIIGaze zhang2015appearance , ➏ TalkingHead-1KH. In addition, we use gray to represent data that is partially uncertain.
@@
HyperReenact bounareli2023hyperreenact 0.710 - 0.230 27.10 -
DG hsu2022dual 0.831 - - 22.10 0.761
AVFR-GAN agarwal2023audio - 32.20 - 08.48 0.824
Free-HeadGAN doukas2023free ➊➌➍➎ 0.810 22.16 0.100 35.40 -
HiDe-NeRF li2023one ➊➋➐ 0.931 21.90 0.084 - 0.862
Table 9: Results of representative face reenactment methods on VoxCeleb for the cross-identity reenactment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ ETH-Xgaze, ➍ Gaze360 kellnhofer2019gaze360 , ➎ MPIIGaze zhang2015appearance , ➏ TalkingHead-1KH.
@@
HyperReenact bounareli2023hyperreenact 0.680 - - - -
AVFR-GAN agarwal2023audio - - - 09.05 -
Free-HeadGAN doukas2023free ➊➌➍➎ 0.789 - - 53.90 13.1
HiDe-NeRF li2023one ➊➋➏ 0.786 0.012 0.971 57.00 -
Table 10: Results of representative face reenactment methods on VoxCeleb2 for quality assessment. Notations: ➊ VoxCeleb, ➋ VoxCeleb2, ➌ LRW, ➍ CelebV-HQ, ➎ TalkingHead-1KH.
@@
PC-AVS zhou2021pose ➋➌ - - 6.880 - 0.886
GC-AVT liang2022expressive - - 2.757 - 0.739
Wang et alwang2023memory - 28.92 1.830 - 0.830
PECHead gao2023high ➋➍➎ 1.590 - 23.05 -
DG hsu2022dual 0.721 - - 51.79 0.540
HiDe-NeRF li2023one ➊➎ 0.787 - - 61.00 -
$\bullet$ Results on Talking Face Generation. Table 11 displays the performance results of various talking face generation approaches on the MEAD wang2020mead dataset since 2023. AMIGO zhai2023talking achieves promising results among all methods, which utilizes seq2seq to generate facial landmarks for emotion tagging, matching mouth movements with emotional features. Additionally, it employs a landmark-to-image translation network to create facial images with fine textures.
$\bullet$ Results on Facial Attribute Editing.  Table 12 evaluates the quality level of generated images using FID, and Table 13 assesses facial reconstruction capabilities using PSNR, LPIPS, and SSIM. Due to different training and testing datasets, quantitative fair comparisons are not possible that just serve as performance demonstrations.
Table 11: Evaluation results of the models involved in talking face generation on the MEAD dataset. Notations: ➊ MEAD, ➋ LRW, ➌ VoxCeleb2, ➍ HDTF.
@@
Xu et alxu2023high 0.83 2.36 - 3.500 15.91 30.09/0.850
EMMN tan2023emmn ➊➋ - - 2.780/2.870 3.570 - 29.38/0.660
AMIGO zhai2023talking ➊➌ - 2.44 2.140/2.440 - 19.59 30.29/0.820
SLIGO sheng2023stochastic 0.88 1.83 - 3.690 - 00000-/0.790
Gan et algan2023efficient ➊➌ - - 2.250/2.470 - 19.69 21.75/0.680
DreamTalk ma2023dreamtalk ➊➌➍ - - 2.910/1.930 3.780 - 00000-/0.860
SPACE gururani2023space ➊➌ - - - 3.610 11.68 0-
TalkCLIP ma2023talkclip - - 3.601/2.415 3.773 - 00000-/0.829
Table 12: FID evaluation of different methods. Notations: ➊ FFHQ, ➋ CelebA-HQ, ➌ MM CelebA-HQ, ➍ CelebA, ➎ CelebAText-HQ, ➏ Self-build.
@@
FENeRF sun2022fenerf GANs+NeRF 12.10
FENeRF sun2022fenerf GANs+NeRF 28.20
AnyFace sun2022anyface GANs 56.75
AnyFace sun2022anyface GANs 50.56
TextFace hou2022textface GANs 22.81
TG-3DFace yu2023towards GANs 52.21
TG-3DFace yu2023towards GANs 39.02
HifaFace gao2021high GANs ➊➋ ➊➋ 04.04
GuidedStyle hou2022guidedstyle GANs 41.79